OpenStudy (anonymous):

can someone help me with this problem ? x/9=8

OpenStudy (debbieg):

To isolate the x, you need to "undo" the division on the left hand side. What operation "undoes" division?

OpenStudy (anonymous):

I am not sure, but it ask me several questions of how to do this.. i have to do a transaction

OpenStudy (debbieg):

If \(\dfrac{a}{b}=c\), then \(b\cdot\dfrac{a}{b}=c\cdot b\), so \(a=cb\)

OpenStudy (debbieg):

(I'm not sure what you mean by "do a transaction".)

OpenStudy (anonymous):

transaction meaning, subtract multiply divide or add both sides

OpenStudy (debbieg):

I won't do your problem for you, but here's a similar example as a reference: \(\Large \dfrac{x}{3}=6\) \(\Large 3\cdot\dfrac{x}{3}=6 \cdot3\) The threes on the left side cancel.... why? because what we REALLY have there is: \(\Large \dfrac{3}{3}\cdot x=6 \cdot3\) \(\Large 1\cdot x=6 \cdot3\) so that means \(\Large x=18\)

OpenStudy (debbieg):

Yes, right - add, subtract, multiply, divide - those are called "operations" to me, but if you are in another part of the world it might be different. :)

OpenStudy (debbieg):

Just think about what the equation is SAYING: it isn't just letters and numbers, it has a meaning! :) For my example, the equation says: x is the number that you divide by 3 in order to get 6. What number is that? It's 18! The algebra just gives us an algorithm for SOLVING it.... but if you think about the mathematical expression in a way that makes intuitive sense, it may help you understand the concepts better.

OpenStudy (anonymous):

lol im in another part of the world alright,, DEF to algebra lol

OpenStudy (anonymous):

ty anyways.. :)

OpenStudy (debbieg):

ok, yw.